OpenStudy (anonymous):

integral from 0 to infinit for 1/x^(1/4) Please help! just wanna check my answer, I got 0

OpenStudy (zzr0ck3r):

\(\int_0^\infty x^{-\frac{1}{4}}dx=\frac{4}{3}x^{\frac{3}{4}}|_0^\infty=\infty\)

OpenStudy (anonymous):

isn't infinity^3/4 = 0

OpenStudy (zzr0ck3r):

no

OpenStudy (zzr0ck3r):

\(\sqrt{\infty}=\infty\)

OpenStudy (dumbcow):

if exponent is negative then infinity^n = 0

OpenStudy (anonymous):

oh ok so then the integral from 1 to infinity would be infinity aswell cause for that I get 4/3infnity^3/4-4/3

OpenStudy (zzr0ck3r):

\(\frac{4}{3}\infty^\frac{3}{4}-\frac{4}{3}*0^\frac{3}{4}=\frac{4}{3}\infty^\frac{3}{4}=\infty\)

OpenStudy (anonymous):

ok thanks!

OpenStudy (zzr0ck3r):

yes

OpenStudy (zzr0ck3r):

I see what you are saying...

OpenStudy (zzr0ck3r):

for any n that is not infinity, negative inifinty

OpenStudy (zzr0ck3r):

you will either subtract or add a finite amount to infinity, so you will get infinity

OpenStudy (zzr0ck3r):

and infinity - infinity does not even make sense.....

OpenStudy (anonymous):

and also I did the integral from 1 to infinityof 1/x^(1/4) if I use p integral rule 1/p-1 I get - 4/3 but if I solve analyticaly and if infinity ^4/3=infinity then I get 0 so that's why im confused